### 04 Jan 2024 v1

# CLAPP: Contrastive Language-Audio Pre-training in Passive underwater vessel classification

Abstract
Existing research on audio classification faces challenges in recognizing attributes of passive underwater vessel scenarios and lacks well-annotated datasets due to data privacy concerns. In this study, we introduce CLAPP (Contrastive Language-Audio Pre-training in Passive Underwater Vessel Classification), a novel model. Our aim is to train a neural network using a wide range of vessel audio and vessel state text pairs obtained from an oceanship dataset. CLAPP is capable of directly learning from raw vessel audio data and, when available, from carefully curated labels, enabling improved recognition of vessel attributes in passive underwater vessel scenarios. Model’s zero-shot capability allows predicting the most relevant vessel state description for a given vessel audio, without directly optimizing for the task. Our approach aims to solve 2 challenges: vessel audio-text classification and passive underwater vessel audio attribute recognition. The proposed method achieves new state-of-the-art results on both Deepship and Shipsear public datasets, with a notable margin of about 7%-13% for accuracy compared to prior methods on zero-shot task.
Index Terms—  Multi-model, Passive underwater vessel audio classification, zero-shot
1 Introduction
The development of precise automatic object classification methods for underwater sensor data, especially Sound Navigation and Ranging (sonar) data, holds significant importance due to its wide range of potential applications [1]. These applications encompass various marine classification tasks, such as inspecting underwater structures for the offshore industry, identifying underwater archaeological remains, monitoring shorelines, quantifying and categorizing marine life behaviors for biological research, and identifying vessels, among others. Additionally, acoustic signals in the marine environment contribute to environmental pollution, impacting marine life and human activities near port areas [2].
Text-to-audio vessel retrieval is a sub-task that falls within the domains of text-audio retrieval and passive underwater audio classification. Different from the traditional image retrieval task [3, 4], passive underwater vessel classification focuses on retrieving a vessel-of-interest from a vast vessel audio database by identifying the best match to a given text description query.
The advancement of underwater acoustic classification techniques has yielded a profound and extensive influence. These techniques have found applications in various domains, resulting in noteworthy advancements.
Illustrative tasks of underwater audio classification encompass applications in the offshore industry for inspecting underwater structures [2, 5], identifying underwater archaeological remains[6], surveilling coastlines [1], quantifying and categorizing marine organism behavior for biological research [7], and discerning ships [8], among various others [9, 10]. Furthermore, acoustic signals within the ocean not only reflect the impacts of environmental pollution on marine organisms [11], but also provide insights into ship activities within nearby harbors [12, 13]. However,
To address the limitations of separate pre-trained models on unimodal datasets, we leverage the Imagebind as the initialization of our model. Imagebind is pre-trained with abundant audio-text pairs and has powerful underlying cross-modal alignment capabilities. With the proposed CLAPP, we successfully transfer the powerful knowledge directly from the pre-trained full Imagebind model and continue to learn fine-grained cross-modal relations on text-to-audio vessel classification datasets. Our main contributions are as follows:
  • We propose the first large-scale, multi-labeled dataset, Oceanship, for pre-training passive underwater vessel audio classification.
  • We redefine intra-domain underwater audio classification tasks and introduce a Domain-Generalization task within this domain. To the best of our knowledge, we are the first to conduct cross-domain experiments in the realm of passive underwater audio classification. Our proposed model, CLAPP, demonstrates a significant 20% improvement in cross-domain effectiveness and achieves performance boosts of 7% to 13% on the Deepship and Shipsear datasets.
  • We leverage the Oceanship dataset to facilitate the development of a multimodal underwater audio classification model. Our work represents a pioneering effort in constructing a pre-trained model specifically tailored for the task of passive vessel audio classification. Additionally, our model achieves state-of-the-art performance through zero-shot learning on the Deepship and Shipsear datasets.
Fig. 1: Overview of the proposed CLAPP framework. It includes a text encoder, an audio encoder, and two Lora adapters arranged in parallel branches. Finally, CLAPP utilizes three different loss functions during the training process: id loss, sdm loss, and sdmrps loss. During the pretraining process of CLAPP, we freeze the parameters of the original audio encoder and text encoder, and then train the parameters of the Lora adapters for corresponding modalities. To achieve the goal of pre-training audio-text classification tasks within limited computational costs.
2 Related work
2.1 Dataset and Benchmark
Shipsear [14] is a database of underwater recordings of ship and boat sounds. It consists of 90 records (nearly 3 hours) that capture sounds from 11 vessel types and one background noise type. DeepShip [15] is an underwater acoustic benchmark dataset raised in 2021. It includes 47 hours and 4 minutes of real-world underwater recordings, encompassing 265 different ships belonging to 4 classes (cargo, passenger ship, tanker, and tug). DeepShip provides category labels for the recordings.
We propose a dataset called oceanship. This dataset comprises 121 hours of genuine underwater recordings, featuring 9,261 distinct vessels categorized into 15 classes. Oceanship is at least 2.5 times larger than existing open-source underwater sound datasets.
Data is the life-blood of training deep neural network models and ensuring their success [16, 17]. As for the passive underwater vessel classification, most of the underwater acoustic models have primarily focused solely on the audio modality [18, 19, 20]. In this context, we specifically address models that exclusively utilize the audio modality and are openly available on open-source datasets. In the field of passive underwater audio classification, UALF [20] draws inspiration from the Leaf approach in conventional audio classification, pioneering the concept of a learnable frontend feature extractor. This approach yields impressive results in terms of classification performance.
2.2 Oceanship Dataset Description
As shown in Table 1, we have compared our dataset with two currently available datasets, DeepShip and Shipsear. In terms of the total duration of the dataset, Oceanship is the largest, featuring 15 categories that encompass 83% of the resolvable vessel names according to the AIS naming protocol. Consequently, the Oceanship dataset encompasses all high-frequency vessels as defined by the AIS naming protocol.
In terms of data volume (duration), the Oceanship dataset is 2.5 times larger than the DeepShip dataset and 70 times larger than the Shipsear dataset. Therefore, the Oceanship dataset offers greater diversity in terms of the data it contains compared to the aforementioned datasets.
Dataset Name Class num Duration(h) Diff in Len(s) AIS Name
Shipsear 11 3.13 <670
Deepship 4 47.22 <300
Oceanship (ours) 15 121.16 <5
Table 1: Comparison of Oceanship with Existing Ship Audio Classification Datasets
3 Task Definition
Intra-Domain Task   The prevailing task can be rigorously delineated within the subsequent context: $S_{train}=\{x_{i},y_{i}\}_{i=0}^{n_{0}}$, $S_{test}=\{x_{j},y_{j}\}_{j=0}^{m_{0}}$; where $\mathcal{P}_{XY}^{train}=\mathcal{P}_{XY}^{test}$. In essence, this formulation implies that both the training and test datasets are derived from the same source. Given that data originating from a single dataset exhibit an identical distribution, the ability of a model to generalize effectively remains unverifiable. To ascertain the model’s capacity for generalization, a prerequisite condition necessitates $\mathcal{Y}_{train}=\mathcal{Y}_{test}$, where $\mathcal{Y}$ delineates the label space. However, the pursuit of generalizability encounters considerable challenges within the paradigm of passive underwater audio classification. Notably, this complexity emerges due to the fundamental distinction between the label spaces of the original datasets, denoted as $\mathcal{Y}_{Deepship}\neq\mathcal{Y}_{Shipsear}$.
[…]
This discernible attribute underscores that the cross-domain framework adopted for evaluation merely encapsulates 40% of the Shipsear dataset and 50% of the Deepship dataset. Consequently, it becomes evident that the utilization of the cross-domain paradigm falls short in harnessing the complete potential of cross-domain data integration.
Zero-shot Task   If the source domain has a particularly small amount of scale data, the performance for domain generalisation will definitely suffer. Here, we define a new task for the underwater audio classification scenario, use a large-scale dataset like Oceanship as the source domain for training, and then use the Deepship and Shipsear datasets for testing during the testing process.
4 Method
In this section, we present our proposed CLAPP framework. The overview of CLAPP is illustrated in Fig. 1, and the details are discussed in the following subsections.
4.1 Feature Extraction Dual-Encoder
Inspired by the partial success of transferring knowledge from CLIP to text-image classification, we directly initialize our CLAPP with the full Imagebind audio and text encoder to enhance its underlying cross-modal alignment capabilities. Furthermore, during the pretraining phase, we employ the Lora adapter to simultaneously train the audio modality branch and the text modality branch while keeping the parameters of the Imagebind backbone network frozen.
Text Encoder. For an input text $T$, we directly use the Imagebind text encoder to extract the text representation, which is a modified version of ViT-Base [21] proposed by Girdhar et al. [22]. Following Imagebind, we employ lower-cased byte pair encoding (BPE) with a vocabulary size of 49152 to tokenize the input text description. The text description is enclosed with [SOS] and [EOS] tokens to indicate the start and end of the sequence. Then, the tokenized text ${f_{\text{sos}}^{t},f_{t}^{1},...,f_{\text{eos}}^{t}}$ is fed into the transformer to exploit correlations between each patch using masked self-attention.
Audio Encoder. We follow the approach described in [22] for encoding audio. We convert a 2-second audio sampled at 16kHz into spectrograms using 128 mel-spectrogram (MFCC) bins. Since the spectrogram is a 2D signal similar to an image, we utilize a ViT model with a patch size of 16 and a stride of 10. Given the spectrogram $A_{\text{spec}}\in\mathbb{R}^{H\times W\times C}$, we employ a pretrained Imagebind ViT model to obtain the image embedding.
To process the spectrogram, we first split it into a sequence of $N=\frac{H\times W}{P^{2}}$ fixed-sized non-overlapping patches, where $P$ represents the patch size. Next, we map the patch sequence to 1D tokens ${f_{i}^{v}}_{i=1}^{N}$ using a trainable linear projection.
LoRA Adapter. LoRA [23] allows us to train some dense layers in a neural network indirectly by optimizing rank decomposition matrices of the dense layers’ change during adaptation instead, while keeping the pre-trained weights frozen as shown in Fig. 1.
4.2 Loss function
During the pre-training phase, we utilize commonly used loss functions for cross-modal alignment, including the ID loss to align different modalities. Additionally, we introduce the SMD loss, which is designed specifically for our framework. Building upon the SMD loss, we propose the SDMRPS loss to address the issue of positive samples appearing in the negative samples when the number of mini-batches is larger than the number of class labels.
ID Loss. The ID loss is a softmax loss which classifies an audio or text into distinct groups based on their identities. It explicitly considers the intra-modal distance and ensures that feature representations of the same audio/text group are closely clustered together in the joint embedding space.
Similarity Distribution Matching. We employed a cross modal matching loss called Similarity Distribution Matching(SDM). Given a mini-batch of N audio-text pairs, for each audio global representation $f^{v}_{i}$ , we construct a set of audio-text representation pairs as $\{(f^{v}_{i},f^{t}_{j}),y_{i,j}\}_{j=1}^{N}$, where $y_{i,j}$ is a true matching label, $y_{i,j}=1$ means that ($f^{v}_{i},f^{t}_{j}$ ) is a matched pair from the same identity, while $y_{i,j}=0$ indicates the unmatched pair. Let $sim(u,v)=\mathbf{u}^{\mathsf{T}}\mathbf{v}/\lVert\mathbf{u}\rVert\lVert\mathbf{v}\rVert$ denotes the dot product between $\mathcal{L}_{2}$ normalized $\mathbf{u}$ and $\mathbf{v}$. Then the probability of matching pairs can be simply calculated with the following softmax function:
$p_{i,j}=\frac{{\exp\left({\text{{sim}}(f^{v}_{i},f^{t}_{j})}/{\tau}\right)}}{{\sum_{k=1}^{N}\exp\left({\text{{sim}}(f^{v}_{i},f^{t}_{k})}/{\tau}\right)}}$
where $\tau$ is a temperature hyperparameter.
Similarity Distribution Matching with Removed multi-positive samples. This module needs to be assisted with the help of mask, that is, Ground Truth module, such as  1, for the contrast learning positive samples (i=j), we mark them as positive samples; for the multiple positive samples labelled as such, we change the value of the position to infinitesimal (-inf), which does not take part in the computation of the contrast learning In loss,the rest of the values are 0 and treated as negative samples;
$v_{ij}=\left\{\begin{array}[]{ll}1&\text{$i=j$}\\ -9e10&\text{$i\neq j$, $v_{ij}=1$}\\ 0&\text{others}\\ \end{array}\right.i,j=0,...,n$ (1)
After that, the objective of using Infonce Loss Equation 2 is to maximise the similarity between the query sample and the positive samples and minimise the similarity between the query sample and the other negative samples. n denotes the number of samples, $z_{i}$ denotes the query sample, $v_{i}$ denotes the positive samples (similar samples that have been stripped of the multiple positive samples labels), and sim($z_{i}$, $v_{j}$) denotes the metric function used to calculate the query similarity between sample $z_{i}$ and positive sample $v_{j}$.
$\mathcal{L}_{\text{a2t}}=-\frac{1}{N}\sum_{i=1}^{N}\log\left(\frac{\exp(\text{sim}(z_{i},v_{i}))}{\sum_{j=1}^{N}\exp(\text{sim}(z_{i},v_{j}))}\right)$ (2)
Symmetrically, the SDMRPS loss from text to audio $\mathcal{L}_{t2i}$ can be formulated by exchanging $z_{i}$ and $v_{i}$ in Eq.1, and the bi-directional SDMRPS loss is calculated by:
$\mathcal{L}_{sdmrps}=\mathcal{L}_{t2a}+\mathcal{L}_{a2t}.$
5 Experiments
In this section, we present our proposed CLAPP framework. The overview of CLAPP is illustrated in Fig. 1 and the details are discussed in the following subsections.
Method pt./ft. dataset task Type Backbone parameters Deepship Acc(%) Shipsear Acc(%)
AudioMAE [24] oceanship zero-shot Audio ViT 86M 26.56 58.45
SSAST [25] oceanship zero-shot Audio ViT 89M 31.93 46.98
LEAF [26] oceanship zero-shot Audio Efficientnet 5.3M 36.18 43.21
Imagebind [22] oceanship zero-shot Audio+Text ViT 180M 38.31 57.23
BEATs [27] oceanship zero-shot Audio Transformer 90M 45.45 62.89
CLAPP(ours) oceanship zero-shot Audio+Text ViT 28.81M 52.35 75.35
UART Shipsear 1% Audio+Text ESResNeXt 50M 55.26 -
CLAPP(ours) Shipsear 1% Audio+Text ViT 28.81M 62.73 -
Table 2: Performance comparisons with state-of-the-art methods on Deepship and Shipsear. Results are ordered based on the accuracy. “Parameters” column stand for trainable parameters in specific model. ”pt.” stand for pre-trained dataset and ”ft.” stand for fine-tuned dataset. ”Audio” and ”Audio+Text” in ”Type” column stand for single modality and multi-modalities.
5.1 Comparison with State-of-the-Art Methods
In this section, we present comparison results with the state-of-the-art models in the field of general audio classification benchmark. Note that the Baseline models in Table 2 refer to various audio classification models fine-tuned with specific loss functions on the Oceanship dataset.
Performance Comparisons on Deepship(zero-shot). We first evaluate the proposed method on the vessel audio classification common benchmark, Deepship. As shown in Tab. 2, CLAPP outperforms all state-of-the-art methods, achieving 52.35% accuracy.
Due to our pioneering work on the zero-shot classification task, current passive underwater classification models do not possess zero-shot classification capability. Therefore, for fair comparison, we pre-trained the state-of-the-art (SOTA) models for general audio classification tasks (such as Audioset, ESC-50, etc.) based on our proposed Oceanship. For the fine-tuning stage, we used the same hyperparameter settings as the CLAPP model and trained for the same number of epochs. In the zero-shot classification task evaluation on Deepship, the CLAPP model outperformed existing multimodal general audio classification SOTA models.
Performance Comparisons on Deepship(1%).
We followed the data classification approach in the UART paper, pre-training on the Shipsear dataset. For transfer learning, we randomly selected 1% of the training data from Deepship and achieved higher results in comparison to existing low-dose data(1%) transfer learning methods in the same field. We experimented with multiple seeds and weighted the results accordingly. Our proposed model achieved a performance of 62.73% on the 1% data, surpassing the current state-of-the-art (SOTA) method in UART.
Performance Comparisons on Shipsear(zero-shot). Our proposed CLAPP model achieved 75.35%, surpassing the recent general audio classification benchmark (BEATs) by +12.5%. As mentioned earlier, all other state-of-the-art (SOTA) general audio classification models were fine-tuned using the same dataset as CLAPP, ensuring a fair comparison.
Evaluation Metrics. We adopt the Accuracy as the primary evaluation metrics. we followed UART to calculate the accuracy according to the official classification criteria (class A–class E) to facilitate comparisons with other works on Shipsear.
Implementation Details. CLAPP consists of a pre-trained audio encoder, i.e., Imagebind-ViT-B/16, a pre-trained text encoder, i.e., Imagebind text Transformer, and Lora Adapter for each trunks. Our model is trained with Adam optimizer for 60 epochs with a learning rate initialized to $1\times 10^{-5}$ and cosine learning rate decay. For random-initialized modules, we set the initial learning rate to $5\times 10^{-6}$. The temperature parameter $\tau$ in SDMRPS loss is set to 0.05. We perform our experiments on a single V100 32GB GPU.
Optimization. As mentioned previously, the main objective of SDM is to improve the learning of global image-text representations in joint embedding space. To achieve this goal, the commonly utilized ID loss is also adopted along with SDM loss and SDMPRS loss to optimize CLAPP. The ID loss is a softmax loss which classifies an audio or text into distinct groups based on their identities. It explicitly considers the intra-modal distance and ensures that feature representations of the same audio/text group are closely clustered together in the joint embedding space. CLAPP is trained in an end-to-end manner and the overall optimization objective for training is defined as:$\mathcal{L}=\mathcal{L}_{id}+\mathcal{L}_{sdm}+\mathcal{L}_{sdmrps}$
No. Loc. Spe. T.S. Deepship Shipsear
0 41.32 69.15
1 50.17 76.43
2 42.63 61.25
3 43.79 63.37
4 50.13 79.27
5 51.45 77.92
6 52.19 71.23
Table 3: For the text modality of CLAPP, we conducted an ablation study on the types of auxiliary information in the input. Here, ”Loc.” represents Location, ”Spe.” represents Speed over ground, and ”TS” represents Timestamp.
No. $\mathcal{L}_{id}$ $\mathcal{L}_{sdm}$ $\mathcal{L}_{sdmrps}$ Deepship Shipsear
0 33.71 67.32
1 37.34 70.12
2 49.12 73.32
3 42.83 74.24
4 51.72 75.12
5 52.38 75.43
6 52.35 75.35
Table 4: The ablation experiments using three types of loss functions on zero-shot task.
5.2 Ablation Study.
In this subsection, we analyze the effectiveness of each component in the CLAPP framework in zero-shot task. Here, we adopt the CLIP-ViT-B/16 model fine-tuned with SDM loss as the Baseline to facilitate the ablation study.
Ablations on auxiliary information. To fully demonstrate the impact of different components in CLAPP, we conduct a comprehensive empirical analysis on two public underwater audio datasets, deepship and shipsear. The accuracy (%) are reported in Tab. 3.
The text modality of our proposed pretraining dataset, Oceanship, encompasses three distinct types of auxiliary information: location information, ship’s speed over ground, and timestamp. By incorporating these auxiliary information into the text modality, our comprehensive analysis from the table demonstrates that the augmentation of auxiliary information positively correlates with improved performance in the zero-shot classification task, specifically on the Deepship and Shipsear datasets.
Ablations on Loss Function. According to the ablation experiments in Tab. 4, the efficacy of IRR is revealed via the experimental results of No.0 vs. No.4 and No.1 vs. No.5. and No.3 vs. No.6. Merely adding the SDMRPS to Baseline improves the accuracy by 18.01%, 15.04%, 9.52% on the two datasets, respectively. The above results clearly show that IRR module are beneficial for cross-modal audio-text matching.
6 Conclusion
In this paper, we introduce a cross modal Contrastive Language-Audio Pre-training in Passive (CLAPP) to learn audio-text representations. Furthermore, we introduce oceanship, which is the first large-scale, multi-labelled scenario and available for pre-training passive underwater vessel audio classification dataset. We further propose a Similarity Distribution Matching with Removed multi-positive samples(SDMRPS) loss to effectively enlarge the variance between non-matching pairs and the correlation between matching pairs. These modules collaborate to align audio and text into a joint embedding space. Significant performance gains on three popular benchmarks datasets prove the superiority and effectiveness of our proposed CLAPP framework. We believe that the ”CLIP-based” approach will be the future trend for underwater vessel audio classification in zero-shot and few-shot task.

### 15 Jan 2024 v2

@@
 No. Loc. Spe. T.S. Deepship Shipsear
-0 41.32 69.15
-1 50.17 76.43
+0 41.32 59.15
+1 50.17 66.43
 2 42.63 61.25
@@
 4 50.13 79.27
-5 51.45 77.92
+5 51.45 67.92
 6 52.19 71.23
@@
 5.2 Ablation Study.
-In this subsection, we analyze the effectiveness of each component in the CLAPP framework in zero-shot task.
+In this subsection, we further analyze the effectiveness of each component in the CLAPP framework in zero-shot task.
 Here, we adopt the CLIP-ViT-B/16 model fine-tuned with SDM loss as the Baseline to facilitate the ablation study.
+More detailed information can be listed as follows:
 Ablations on auxiliary information.
-To fully demonstrate the impact of different components in CLAPP, we conduct a comprehensive empirical analysis on two public underwater audio datasets, deepship and shipsear.
-The accuracy (%) are reported in Tab.
+To fully demonstrate the impact of different components in CLAPP, we conduct a comprehensive empirical analysis on two public underwater audio datasets, named deepship and shipsear respectively.
+More detailed accuracy (%) performance are reported in Tab.
 3.
-The text modality of our proposed pretraining dataset, Oceanship, encompasses three distinct types of auxiliary information: location information, ship’s speed over ground, and timestamp.
-By incorporating these auxiliary information into the text modality, our comprehensive analysis from the table demonstrates that the augmentation of auxiliary information positively correlates with improved performance in the zero-shot classification task, specifically on the Deepship and Shipsear datasets.
+In general, the text modality of our proposed pretraining dataset, Oceanship, encompasses three distinct types of auxiliary information: location information, ship’s speed over ground, and timestamp.
+According to the experimental results in Tab.
+3, by incorporating these auxiliary information into the text modality, our comprehensive analysis from the table demonstrates that the augmentation of auxiliary information positively correlates with improved performance in the zero-shot classification task, specifically on the Deepship and Shipsear datasets.
 Ablations on Loss Function.